# JointDiT: Enhancing RGB-Depth Joint Modeling with Diffusion Transformers

### v2

@@ 5 Related Work
The success of DDPM [23] has demonstrated the effectiveness of diffusion models for text-to-image generation, utilizing a forward and backward process formulated as a Markov chain. Score-based generative models [50, 51] provided a different perspective by modeling the diffusion process as learning the score, i.e., the gradient of the log probability density, from noisy data. This approach was further extended with Stochastic Differential Equations (SDEs), which unify the forward and backward processes in a continuous-time framework [52]. More recently, Flow Matching [35] has been introduced as an alternative to diffusion models, enabling exact likelihood training through Continuous Normalizing Flows (CNFs) [8].
@@
Text-to-image diffusion models have demonstrated their effectiveness in conditional and joint generation tasks. ControlNet [61] introduced an additional zero-initialized network, enabling fine-grained control over conditional generation tasks such as depth-to-image and pose-to-image synthesis. Building on this, LooseControl [4] proposed a more relaxed conditioning approach, allowing for weaker or more flexible integration of conditional information, thereby broadening the applicability of diffusion models in various generative tasks.
Meanwhile, prior works [28, 18, 19] have shown that the image prior of pre-trained diffusion models can be beneficial for depth estimation. Models such as JointNet [60], LDM3D [53], and UniCon [31] were designed to model the joint distribution between images and depth maps using diffusion models. However, these models are based on a U-Net based diffusion architecture, which has a limited receptive field. This is in contrast to recent findings suggesting that diffusion transformers provide a stronger image prior and a global receptive field, which is particularly useful for dense prediction tasks [43, 33, 1]. In this paper, we leverage the advantages of diffusion transformer to model the solid joint distribution between image and depth.
@@ 6 Conclusion
We propose JointDiT, which models solid joint distribution. By harnessing the strong image prior and global receptive property of the state-of-the-art diffusion transformer, we build a unified model capturing multi-modal joint distribution at any separate noise levels. To achieve solid distribution, we present two simple yet effective techniques, called the adaptive scheduling weights dependent on the noise levels of modalities and unbalanced timestep sampling strategy. Through comprehensive experiments, we show that these two techniques notably improve the performance of joint generation, depth estimation, and depth-to-image generation. Our complete model generates images and depth maps, which form highly plausible and image-aligned 3D structures when lifted into 3D space. Furthermore, JointDiT achieves depth estimation performance comparable to that of diffusion-based depth estimation models, demonstrating that a joint distribution model can serve as a replacement for a conditional distribution model.
@@ B    Additional Experiments
  B.1    Jointly Generated Image Quality
  B.2    Ablation of the LoRA’s Rank
  B.3    Joint Panorama Generation
@@ Appendix A Implementation Details
We provide the details of experiment setup, dataset preprocessing, and our method.
@@ A.1 Experiment Setup
We will describe in detail the configurations we used for joint generation, depth estimation, and depth-conditioned image generation. We consistently use 20 denoising steps and a guidance scale of 3.5 across all experiments.
@@
We generate images and their corresponding depth maps by initially setting $t_{x}=0$ and $t_{y}=0$, i.e., sampling from a standard normal distribution. In the main paper, all joint generation experiments are conducted using text prompts. However, we observe that joint generation occurs even without a text prompt. To compare with JointNet [60] and LDM3D [53], we generate 512×512 images and depth maps jointly. Despite being trained only on a 512×512 resolution dataset, we observe that JointDiT successfully operates at varying resolutions, such as 1024×1024.
@@
Table 5: LoRA-applied components. To build the depth branch extending the original Flux model [5], we add LoRAs to MM-DiT, P-DiT, and Input stage.
@@
To estimate the depth map from a given image, we set $t_{x}=1$ and $t_{y}=0$ and provide an empty text prompt. We do not use the ensemble technique differently from Marigold [28] and Geowizard [18]. Since JointDiT can operate at varying resolutions, we use the NYUv2, ScanNet, KITTI, and DIODE datasets [49, 14, 3, 55] at their original resolutions as model inputs. For the ETH3D dataset [48], which has a 4K resolution, we resize the images while preserving the aspect ratio so that the larger dimension is set to 1024 pixels.
@@
We generate depth-conditioned images from given text prompts by initially setting $t_{x}=0$ and $t_{y}=1$. The conditioning depth maps are obtained by Depth-Anything-V2 [59].
@@ A.2 Data Preprocessing
Generation Method ImageNet 6K Pexel 6K MSCOCO 30K
@@ A.3 Unbalanced Timestep Sampling Strategy
When applying the unbalanced timestep sampling strategy, the timesteps, i.e., $t_{x}$ and $t_{y}$, are separately sampled from the timestep distributions $f(t)$ and $g(t)$, respectively, or vice versa. This is applied with a 50% probability during training, while for the remaining 50%, the same timestep sampled from $f(x)$ is used for both $t_{x}$ and $t_{y}$. The timestep distribution is as follows:
$f(t)=\frac{\sigma(z)\cdot s}{1+(s-1)\cdot\sigma(z)},\quad\text{where }z\sim\mathcal{N}(0,1).$ (7)
@@
A.4 Architecture of Depth Branch
To build the depth branch, we add LoRAs [24] to the original Flux architecture [5]. We added LoRA to the components connected before and after the attention mechanisms of the multi-modal diffusion transformer (MM-DiT) and parallel diffusion transformer (P-DiT) blocks [16, 15] that constitute Flux. Table 5 demonstrates the LoRA applied components of MM-DiT and P-DiT blocks. We additionally add LoRAs to the input stage.
@@ Appendix B Additional Experiments
B.1 Jointly Generated Image Quality
Method ImageNet 6K Pexel 6K MSCOCO 30K
@@ B.1 Jointly Generated Image Quality
Rank1$\uparrow$ Rank2 Rank3$\downarrow$ Rank1$\uparrow$ Rank2 Rank3$\downarrow$ Rank1$\uparrow$ Rank2 Rank3$\downarrow$
@@
Table 7: Human preference evaluation on images jointly generated by joint generation methods [53, 60] and Ours. We access the human preference by utilizing ImageReward [58] that was trained to estimate human preference. With both joint generation models and ours, we conduct joint generation using the same text prompts and rank the results with ImageReward, obtaining the percentage for each ranking. Our JointDiT achieved the highest rank 1 percentage and the lowest rank 3 percentage across all datasets.
We quantitatively compare the quality of jointly generated images from JointNet [60], LDM3D [53], and our method. For evaluation, we use the dataset from Section 4.4 of the main paper, i.e., ImageNet 6K, Pexel 6K, and MSCOCO 30K. We measure the Inception Score (IS) [47], Fréchet Inception Distance (FID) [21], and CLIP similarity [41] as our evaluation metrics. Table 6 summarizes the results. We also include the results of baseline diffusion models, i.e., Stable diffusion [45] and Flux [5]. Interestingly, Flux achieves relatively high FID scores across all evaluation datasets despite its outstanding text-to-image generation capability. We observe that Flux often generates stylized images. Figure 6 shows samples from ImageNet 6K and the corresponding images generated by Flux. The generated samples appear surreal, which leads to a higher FID between them and the real image dataset. Our model achieves a lower FID score than Flux by learning the joint distribution of images and their corresponding depth maps on the real dataset. However, our IS score is lower than that of Flux, likely due to the limited size of the training dataset.
@@
B.2 Ablation of the LoRA’s Rank
We adopted a LoRA [24] rank of 64 in our JointDiT model. To analyze the effect of the LoRA rank, we train our model with different LoRA ranks and evaluate depth estimation performance on the NYUv2 and ScanNet datasets [49, 14]. As shown in Table 8, as the LoRA rank increases, the depth estimation performance improves, achieving the best performance at the LoRA rank of 64. We did not increase the LoRA rank beyond 64 because the number of model parameters grows exponentially.
B.3 Joint Panorama Generation
JointDiT can be used for RGB-D panorama generation as well. For panorama generation, we strictly follow to the JointNet [60] method combining whole and tile-based denoising strategies [2, 25], to ensure a fair comparison. We denoise image and depth tiles by using joint generative diffusion models. During only early steps, we perform denoising on the entire panorama, and throughout all steps, we aggregate model estimations from both overlapped individual tiles and the whole panorama. Figure 7 demonstrates the RGB-D panorama results. Compared to JointNet, JointDiT shows clear and structurally reasonable images along with sharp depth maps.
@@ Appendix C Additional Qualitative Results
AbsRel $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ $\delta_{1}$uparrow
@@
Utilizing our JointDiT model, we generate images and corresponding depth maps. We visualize the images and depths with their 3D lifting results. As shown in Figure 8, our joint generation results are geometrically reasonable in 3D, with the surface characteristics of the images being well-preserved in the 3D space (e.g., smooth or rough textures).

### v3

@@ 5 Related Work
The success of DDPM [26] has demonstrated the effectiveness of diffusion models for text-to-image generation, utilizing a forward and backward process formulated as a Markov chain. Score-based generative models [57, 58] provided another perspective by modeling the diffusion process as learning the score, i.e., the gradient of the log probability density, from noisy data. It was further extended with Stochastic Differential Equations (SDEs), which unify the forward and backward processes in a continuous-time framework [59]. More recently, Flow Matching [39] has been introduced as an alternative to diffusion models, enabling exact likelihood training through Continuous Normalizing Flows (CNFs) [11].
@@
Text-to-image diffusion models have demonstrated their effectiveness in conditional and joint generation tasks. ControlNet [72] introduced an additional zero-initialized network, enabling fine-grained control over conditional generation tasks such as depth-to-image and pose-to-image synthesis. Building on this, LooseControl [6] proposed a more relaxed conditioning approach, allowing for weaker or more flexible integration of conditional information. Meanwhile, the image prior of pre-trained diffusion models can be beneficial in a wide range of computer vision tasks, such as sound-to-image generation [62, 61], single-image 3D reconstruction [41, 70], and 3D object texturing [9, 69].
In depth-related tasks, prior works [32, 21, 22] have demonstrated the effectiveness of diffusion priors for depth estimation. Hyoseok et al[28] further showed that such priors can be used to solve inverse problems, specifically depth completion. Models such as JointNet [71], LDM3D [60], and UniCon [35] were designed to model the joint distribution between images and depth maps using diffusion models. However, these models are based on a U-Net based diffusion architecture, which has a limited receptive field. This is in contrast to recent findings suggesting that diffusion transformers provide a stronger image prior and a global receptive field, which is particularly useful for dense prediction tasks [50, 37, 1]. In this paper, we leverage the advantages of diffusion transformer to model the solid joint distribution between image and depth.
@@ 6 Conclusion
We propose JointDiT, which models solid joint distribution. By harnessing the strong image prior and global receptive property of a state-of-the-art diffusion transformer, we build a unified model capturing multi-modal joint distribution at any separate noise levels. To achieve solid distribution, we present two simple yet effective techniques, called the adaptive scheduling weights dependent on the noise levels of modalities and unbalanced timestep sampling strategy. Through comprehensive experiments, we show that these two techniques notably improve the performance of joint generation, depth estimation, and depth-to-image generation. Our complete model generates images and depth maps, which form highly plausible and image-aligned 3D structures when lifted into 3D space. Furthermore, JointDiT achieves depth estimation performance comparable to that of diffusion-based depth estimation models, demonstrating that a joint distribution model can serve as a viable alternative for conditional distribution models.
@@
Acknowledgment
This work was supported by the MSIT (Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program (RS-2024-00436680) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation). It was also supported by the KAIST cross-generation collaborative lab project, the Ministry of Science and ICT and NIPA (HPC Support Project), and Microsoft Research Asia.
@@ B    Additional Experiments
  B.1    Advantages of Joint RGB-Depth Modeling
  B.2    Jointly Generated Image Quality
  B.3    Ablation of the LoRA’s Rank
  B.4    Analysis of Failure Cases
  B.5    Joint Panorama Generation
@@ Appendix A Implementation Details
We provide the details of the experiment setup, dataset preprocessing, proposed unbalanced timestep sampling strategy, and architecture design of JointDiT.
@@ A.1 Experiment Setup
We will describe in detail the configurations we used for joint generation, depth estimation, depth-conditioned image generation, and Joint RGB-Depth feature visualization. We consistently use 20 denoising steps across all experiments.
@@
We generate images and their corresponding depth maps by initially setting $t_{x}=0$ and $t_{y}=0$ by sampling noises from a standard normal distribution. While the main paper presents joint generation results conditioned on text prompts, we find that joint generation occurs even without a text prompt. To compare with JointNet [71] and LDM3D [60], we generate 512×512 images and depth maps jointly. Despite being trained only on a 512×512 resolution dataset, we observe that JointDiT successfully operates at varying resolutions, such as 1024×1024.
@@
Table 5: LoRA-applied components. To build the depth branch extending the original Flux model [7], we add LoRAs to $M$-DiT, P-DiT, and Input stage.
@@
To estimate the depth map from a given image, we set $t_{x}=1$ and $t_{y}=0$ and provide an empty text prompt. Unlike Marigold [32] and Geowizard [21], we do not use any ensemble technique. Since JointDiT can operate at varying resolutions, we use the NYUv2, ScanNet, KITTI, and DIODE datasets [56, 17, 5, 64] at their original resolutions as model inputs. For the ETH3D dataset [55], which has a 4K resolution, we resize the images while preserving the aspect ratio so that the larger dimension is set to 1024 pixels. This preprocessing strategy is consistently applied to the comparison methods as well, and for methods that require a fixed input resolution, we use their designated resolution for evaluation.
@@
We generate depth-conditioned images from given text prompts by initially setting $t_{x}=0$ and $t_{y}=1$. The conditioning depth maps are obtained by Depth-Anything-V2 [68]. For the experiment of Sec. 4.3 in the main paper, we follow the evaluation setting of UniCon [35] to compare with Readout-Guidance [45], ControlNet [72], and UniCon. Specifically, we train our model and these methods on the same training dataset, which includes 16k images of PascalVOC [20], depth maps from Depth-Anything-V2 [68], and text prompts extracted using BLIP2 [34]. For the evaluation, using the selected 6k images from the OpenImages dataset [33], we estimate depth maps using an off-the-shelf model and generate images conditioned on these depth maps and text prompts from BLIP2.
Joint RGB-Depth feature visualization
For the feature visualization of Sec. 4.4 in the main paper, we strictly follow the method proposed by Tumanyan et al[63], and visualize the PCA results of the features from each MM-DiT block. Similar to Tumanyan et al., who collected images from semantically related domains (such as humanoid pictures) for visualization, we perform joint generation on 50 samples for each domain, i.e., pixel art style illustrations and indoor scenes that are used in the two examples shown in Fig. 6 of the main paper. We extract features at approximately 50% of the generation process (i.e., $t=0.48$), and apply PCA to visualize them. Due to the architecture structure of the Flux model, which applies positional encoding immediately before every attention layer, we subsample the even indices before applying PCA.
@@ A.2 Data Preprocessing
Figure 7: Comparison of 3D lifting results from our JointDiT, Marigold, and Depth-Anything-V2. The jointly generated depth from JointDiT leads to more coherent 3D shapes and better preservation of structural details compared to the estimated depths.
Generation Method ImageNet 6K Pexels 6K MSCOCO 30K
@@ A.3 Unbalanced Timestep Sampling Strategy
When applying the unbalanced timestep sampling strategy, the timesteps, i.e., $t_{x}$ and $t_{y}$, are separately sampled from the timestep distributions $f(t)$ and $g(t)$, respectively, or vice versa. This is applied with a 50% probability during training, while for the remaining 50%, the same timestep sampled from $f(t)$ is used for both $t_{x}$ and $t_{y}$. The timestep distribution is as follows:
@@
[…]
A.4 Architecture of JointDiT
To build the depth branch, we add LoRAs [27] to the original Flux architecture [7]. Specifically, we add LoRAs to the components connected before and after the attention mechanisms of the multi-modal diffusion transformer (MM-DiT) and parallel diffusion transformer (P-DiT) blocks [19, 18] that constitute Flux. Table 5 summarizes the LoRA-applied components in the MM-DiT and P-DiT blocks. We use a LoRA rank of 64 for both MM-DiT and P-DiT, and apply relatively larger ranks of 512 or 1024 to the input stage. The alpha value is set to half of the corresponding rank.
To design the joint connection module, we adopt the joint cross-attention module from UniCon [35], followed by a zero-initialized linear projection layer. The adaptive scheduling weight is applied subsequently.
@@ Appendix B Additional Experiments
B.1 Advantages of Joint RGB-Depth Modeling
As mentioned in the main paper, we observe that joint RGB-Depth generation tends to yield more plausible 3D lifting results compared to estimating depth from generated images. Figure 7 presents the 3D lifting results by showing top and side views. When using the depth generated by our JointDiT, the results exhibit more well-structured and volumetric 3D geometry than those produced by Marigold [32] and Depth-Anything-V2 [68].
Furthermore, as also discussed in the main paper, our joint generation approach enables plausible depth synthesis even in illustration domains, where depth estimation methods often struggle. Additional qualitative results are presented in Figure 12.
B.2 Jointly Generated Image Quality
Method ImageNet 6K Pexels 6K MSCOCO 30K
@@ B.2 Jointly Generated Image Quality
[…]
@@
Table 7: Human preference evaluation on images jointly generated by joint generation methods [60, 71] and Ours. We assess the human preference using ImageReward [67] that was trained to estimate human preference. With both joint generation models and ours, we conduct joint generation using the same text prompts and rank the results with ImageReward, obtaining the percentage for each ranking. Our JointDiT achieved the highest rank 1 percentage and the lowest rank 3 percentage across all datasets.
We quantitatively compare the quality of jointly generated images from JointNet [71], LDM3D [60], and our method. For evaluation, we use the dataset from Section 4.4 of the main paper, i.e., ImageNet 6K, Pexels 6K, and MSCOCO 30K. We measure the Inception Score (IS) [54], Fréchet Inception Distance (FID) [24], and CLIP similarity [48] as our evaluation metrics. Table 6 summarizes the results. We also include the results of baseline diffusion models, i.e., Stable diffusion [52] and Flux [7]. Interestingly, Flux achieves relatively high FID scores across all evaluation datasets despite its outstanding text-to-image generation capability. We observe that Flux often generates stylized images. Figure 8 shows samples from ImageNet 6K and the corresponding images generated by Flux. The generated samples appear surreal, which leads to a higher FID between them and the real image dataset. Our model achieves a lower FID score than Flux by learning the joint distribution of images and their corresponding depth maps on the real dataset. However, our IS score is lower than that of Flux, likely due to the limited size of the training dataset.
@@
B.3 Ablation of the LoRA’s Rank
We adopt a LoRA rank of 64 in the DiT blocks of our JointDiT model. To analyze the effect of the LoRA rank, we train our model with different LoRA ranks and evaluate depth estimation performance on the NYUv2 and ScanNet datasets [56, 17]. As shown in Table 8, as the LoRA rank increases, the depth estimation performance improves, achieving the best performance at the LoRA rank of 64. We did not increase the LoRA rank beyond 64 because the number of model parameters grows exponentially.
B.4 Analysis of Failure Cases
We observe that our method shares similar limitations with depth estimation methods [68, 32], particularly in handling reflective surfaces such as mirrors. As shown in Fig. 9, both our model and depth estimation models fail to recognize mirrors as flat and planar regions.
B.5 Joint Panorama Generation
JointDiT can be used for RGB-D panorama generation as well. For panorama generation, we strictly follow the JointNet [71] method combining whole and tile-based denoising strategies [4, 29], to ensure a fair comparison. We denoise image and depth tiles by using joint generative diffusion models. During only early steps, we perform denoising on the entire panorama, and throughout all steps, we aggregate model estimations from both overlapped individual tiles and the whole panorama. Figure 10 demonstrates the RGB-D panorama results. Compared to JointNet, JointDiT shows clear and structurally reasonable images along with sharp depth maps.
@@ Appendix C Additional Qualitative Results
AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$
@@
Figure 9: Failure cases in depth estimation. Red and Blue areas indicate near and far depth predictions, respectively.
@@
Utilizing our JointDiT model, we generate images and corresponding depth maps. We visualize the images and depths with their 3D lifting results. As shown in Fig. 11, our joint generation results are geometrically reasonable in 3D, with the surface characteristics of the images being well-preserved in the 3D space (e.g., smooth or rough textures). Furthermore, Figure 12 highlights the effectiveness of our joint generation approach in illustration domains, where plausible 3D structures are obtained despite the inherent difficulty of estimating geometry from stylized images.
@@
Figure 12: Joint generation results in illustration domains. The jointly generated images and depths from JointDiT produce geometrically plausible 3D structures, even in stylized domains.